xdy/dx–y=e^(dy/dx)
1
Expert's answer
2020-02-13T10:02:59-0500

"x\frac{dy}{dx}-y=e^{\frac{dy}{dx}}"

Let's rewrite the equation in form

"y=xy'-e^{y'}"

It's a Clairault's equation.

Differentiation with respect to x will yield

"y'=y'+xy''+e^{y'}y''"

Therefore

"y''\left(x-e^{y'}\right)=0"

Hence, either

"y''=0\iff y=c_1x+c_2" - which is the general solution of Clairault's equation.

or

"x-e^{y'}=0"

"\ln x=y'"

"y=x\ln x-x+c" - which is the singular solution of Clairault's equation.